find the vertex, the focus, and directrix. y^2= 1/2x
Answer by lwsshak3(11628)

find the vertex, the focus, and directrix. y^2= 1/2x
**
This is an equation of a parabola that opens rightwards of the standard form:
(y-k)^2=4p(x-h), (h,k)=(x,y) coordinates of the vertex.
For given parabola:
vertex:(0,0)
axis of symmetry: x-axis or y=0
4p=1/2
p=1/8
focus: (1/8,0) (p distance from right of vertex on axis of symmetry)
directrix: x=-1/8 (p distance from left of vertex on axis of symmetry)